OpenStudy (anonymous):

A 75.0-ml volume of 0.200M NH3 (Kb=1.8 x 10^-5) is titrated with 0.500M NHO3 . Calculate the pH after the addition of 13.0mL of NH3.

OpenStudy (anonymous):

This forms a buffer. meq of NH3=75*0.2=15 meq of HNO3=0.5*13=6.5 meq of NH3 left=8.5. meq of NH4+ = 6.5 By Henderson's equation pOH= pKb + log (base/salt) pOH=5-log1.8+log85/65 pOH=4.86 pH=14-4.86=9.14